OpenStudy (anonymous):

Find the magnitude ||v|| and direction angle ), to the nearest tenth of a degree, for the given vector v. v= -4i - 3j I have no clue how to work this any help?

OpenStudy (perl):

Let vx and vy be the x and y components of v. magnitude v = sqrt( vx^2 + vy^2 ) = sqrt( (-4)^2 + (-3)^2 )